# Generalized Compressed Sensing for Image Reconstruction with Diffusion Probabilistic Models

### v3

@@ 2.1 Linear inverse problem
Given an image $x\in\mathcal{R}^{d}$, we express a linear measurement as $m=M_{k}^{T}x$, where $M_{k}\in\mathcal{R}^{d\times k}$, is a measurement matrix, and $m\in\mathcal{R}^{k}$ is the measurement which provides a partial observation of $x$ (i.e., $M_{k}$ is low rank, $k<d$), We assume $m$ noise-free. The linear inverse problem is to reconstruct an approximation of the original image from the measurement, $\hat{x}(m)$, where $\hat{x}(\cdot)$ can be nonlinear.
@@ 2.2 Image prior embedded in a denoiser
where $y$ is the noise-corrupted signal: $y=x+z$, $z\sim\mathcal{N}(0,\sigma^{2}\text{I})$, and $\hat{x}(y)$ is the MMSE denoising solution. This remarkable equation provides an explicit connection between denoising and the density of the noisy image $p_{\sigma}(y)$. See Appendix A for the proof of Equation 1. The distribution of noisy images, $p_{\sigma}(y)$ is related to the image prior $p(x)$ through marginalization:
[…]
where $g_{\sigma}(z)$ is the distribution of Gaussian noise with variance $\sigma^{2}$. This equation expresses the convolution of $p(x)$ with the Gaussian probability density function. That is, $p_{\sigma}(y)$ is a blurred version of $p(x)$ where the extent of blur depends on $\sigma$. The family of $p_{\sigma}(y)$ over all $\sigma>0$ forms a scale-space representation of $p(x)$ and is akin to the temporal evolution of a diffusion process of $p(x)$. A learned denoiser trained over a wide range of $\sigma$ approximates this family of gradients of log densities and can then be used in a coarse-to-fine gradient ascent algorithm to sample from $p(x)$ (see Algorithm 1).
@@ 2.3 Inverse problem as constrained sampling
See Algorithm 2 for a detailed description of the constrained sampling algorithm.
@@
which can be approximated by averaging over multiple conditional samples for a given measurement model $M_{k}$, see Section E.3 for more details. Visually, however, individual samples will look sharper and of higher visual quality compared to the MMSE estimate (Kadkhodaie & Simoncelli, 2021; Kawar et al., 2022).
@@ 2.4 Optimized linear measurement
In order to solve the optimization problem of Equation 6, we use stochastic gradient descent in the space of orthonormal matrices. Specifically, we parameterize the set of matrices $Q\in\mathcal{R}^{d\times k}$ with orthonormal columns using the Householder product, which represents matrices as a sequence of elementary reflections as the following (Trefethen & Bau, 2022; Shepard et al., 2015):
@@
Here, each elementary reflector $H_{i}$ defines a reflection around a plane. Each vector $v_{i}$ is of the form $[\mathbf{0},1,u^{i}]^{T}$, with the first $i-1$ elements being zero, and $\tau_{i}$ is a scale factor: $\tau_{i}=2/(1+||v_{i}||^{2})$. The collection of $u_{i}$’s forms a lower triangular matrix of $\mathcal{R}^{d\times k}$. They are the free parameters $\phi$ of the parameterization $Q(\phi)$. We can thus search for measurement matrices in the space of $\phi$.
We re-write our empirical objective function from Equation 5 using the parameterization $Q(\phi)$:
$\mathcal{L}(\phi)=\frac{1}{N}\sum_{i=1}^{N}||h(Q(\phi)^{T}x_{i};\ Q(\phi))-x_{i}||_{2}^{2}.$ (8)
We search for the optimal measurement matrix within $Q(\phi)$ through variants of gradient descent (Adam optimizer (Kingma & Ba, 2014), see Supplementary Appendix E):
@@
In practice, the gradient $\nabla_{\phi_{t}}\mathcal{L}(\phi_{t})$ is approximated by computing the MSE (Equation 8) on a subset of images sampled from the training set on each iteration. The gradient descent formulation is general, and applicable to any differentiable objective. As a demonstration of this, we also explore a perceptual loss function, the structural similarity index measure (SSIM) (Wang et al., 2004), using the implementation in Detlefsen et al. (2022). See Appendix E for details of the datasets, network training, and linear measurement optimization.
@@ 3.1 Two-dimensional example
Figure 1: Reconstruction of 2-D signals from a 1-D linear measurement. Top row: Gaussian distribution. Middle row: K-Sparse distribution. Bottom row: A data distribution confined to a curved one-dimensional manifold. A) Samples from the target distribution (blue points). Measurements are obtained via dot product between samples and a unit vector (black vector) parameterized by angle $\theta$. B) Mean squared error (MSE) of sample reconstruction as a function of measurement angle $\theta$, for two estimators: linear projection onto the measurement vector (blue), and reconstruction using the denoiser prior (orange). The dashed vertical line indicates the optimal solution, $\theta^{*}$, obtained with our optimization procedure (based on Equation 8). C) Test data drawn from the prior (blue), plotted in a new coordinate system $(x_{1}^{*},x_{2}^{*})$ rotated by $-\theta^{*}$ (i.e., such that the optimal measurement $m$ corresponds to the horizontal coordinate $x_{1}^{*}$). Example bivariate estimates sampled from the denoiser prior conditioned on the optimal linear measurement are plotted in red. The MMSE solution (orange) is obtained by averaging over multiple conditional samples.
@@
As an illustration of our method, consider a 1-D measurement problem in a 2-D signal space, for three different prior probabilities (Figure 1A): a correlated Gaussian distribution, a K-sparse (union of subspace) model, and a tight distribution along a closed 1D manifold. In each case, we train a small, two-layer fully connected denoising network on the data distribution, which is then used to solve the inverse problem via Algorithm 2. The mean squared reconstruction error obtained by this method as a function of the measurement vector angle is shown in Figure 1B (red). As a baseline, we compare this to the reconstruction obtained by linear projection onto the measurement axis (Figure 1B, blue). The optimal measurement vector is obtained by evaluating the reconstruction error for a set of densely sampled 2-D unit measurement vectors spanning orientations $\theta\in[-\pi/2,\pi/2]$ (Figure 1B, orange). For the optimal measurement vector, Figure 1C shows samples conditioned on each measurement value (red points) and the MMSE solution as a function of the measurement value (orange line).
@@
We first consider a bivariate Gaussian distribution (Figure 1, top row). In this case, the first PC is the optimal measurement for both reconstruction methods, as expected for a Gaussian prior. The second row of Figure 1 depicts the result for a union-of-subspace sparse distribution. Here, the first PC is aligned with the horizontal axis. However, the optimal measurement vector for the denoiser prior is dramatically different. The reconstruction error at the optimal $\theta^{*}$ is near zero. This recapitulates the classical compressed sensing result (Donoho, 2006; Tropp, 2006) that for sparse signals, improved estimates can be achieved by making off-axis measurements.
@@
The last row illustrates the scenario of primary interest in this paper: The data distribution lies in a low-dimensional but curved manifold. This type of higher-order structure cannot be adequately captured by either the Gaussian or sparse prior, but can be effectively described using the more powerful diffusion models, such as our denoiser prior (Supplemental Figure 2). Notably, in this case, we also identified an optimal measurement angle that outperforms the principal axis. Importantly, the optimal angle $\theta^{*}$ is correctly identified by our optimization method across all three cases, illustrating the generality of our approach.
@@
Finally, we highlight another aspect of the reconstruction demonstrated by the 2D examples (Figure 1C). Our methods generate estimates by sampling from the denoiser prior, conditioned on the linear measurements. While individual samples are consistent with the conditional prior, the MMSE solutions are obtained by averaging multiple conditional samples. This averaging, however, can cause deviations of the MMSE solution from the prior. In the context of image reconstruction, the MMSE solution often exhibits lower visual quality (e.g., blurrier images) compared to individual conditional samples, as it does not lie close to the image prior manifold (Kadkhodaie & Simoncelli, 2021; Kawar et al., 2022). This observation is particularly relevant in the results presented below and further motivates our approach of optimizing measurements with respect to a perceptual loss function later.
@@ 3.2 Optimized measurements for MNIST
We used our method to obtain optimized measurements for the MNIST dataset (Deng, 2012). We first trained a DNN denoiser on these digit images – for details of the architecture and training, see Appendix E. We apply our method as described in Section 2.4 to obtain the optimized measurement matrix (OLM) for this dataset, for a range of $k$ values.
@@
For comparison, we apply our reconstruction method to two other types of linear measurements: the top $k$ PCs, and $k$ random vectors, which are optimal measurements under Gaussian and sparse prior assumptions, respectively. We also compare to the optimal reconstruction for a Gaussian prior, which is simply linear projection onto the space spanned by the top $k$ PCs. The top half of Figure 2 illustrates our results for a test digit image. First, we observe that combining the PCs with the denoiser prior significantly improves the linear inverse estimates, which reflects the power of the denoiser prior. All conditional samples from the denoiser prior appear to be real MNIST images, which then gradually converge to the original image as $k$ increases. Use of the optimal measurement matrix (OLM) offers additional improvements, reconstructing the correct digit with a smaller $k$. These improvements are evident in both the identity of the digits and their more detailed appearance. Using random measurements (not shown) does not perform as well as either the optimized linear measurements or PCA. See Supplemental Figure 3 for two more examples, including random measurement vectors.
@@
Figure 3: Performance comparisons for A) MNIST and B) CelebA datasets. Left graph in each panel shows the peak signal-to-noise ratio (PSNR, in dB) as a function of the number of measurements $k$, for the denoiser prior reconstruction from OLMs (red), denoiser prior reconstruction from PCs (orange), linear (projection) reconstruction from PCs (blue), and denoiser prior reconstruction from random measurements (green). Inline numbers indicate the increase in PSNR from PC + denoiser prior to OLM + prior. The right graph in each panel plots the PSNR of the PC reconstruction against that of the OLM reconstruction over all images in the test set, for a value of $k$ approximately $10\%$ of the total number of pixels. Red points indicate the examples shown in Figure 2 and Supplemental Figure 3 and Figure 4. Also see Supplemental Table 1 for comparison of our results to other previous methods.
@@
We quantify our results in terms of peak signal-to-noise ratio (PSNR), as a function of $k$ in the left panel of Figure 3A. The OLMs result in superior reconstruction (red curve), particularly for the values of $k$ ranging from $1\%$ to $30\%$ of the total number of pixels. The peak performance gain of OLM over PC (both utilizing denoiser prior) reached 4.72 dB in PSNR at $k=256$. Between the two PC reconstructions, the nonlinear prior-based method (orange) shows a considerable improvement over linear reconstruction (blue), demonstrating the substantial advantage of using the more complex denoiser prior. See Supplemental Table 1 for a comparison of our results to other methods in the literature. Lastly, consistent with previous reports using simpler image priors (Weiss et al., 2007), we observe that PC measurements outperform random measurements when used with the denoiser prior (green). This raises the possibility that the union of subspace prior assumed in the compressed sensing literature (Donoho, 2006) does not fully capture the structure of these images, since random measurements are expected to perform well if the images indeed lie on a union of subspaces – a structure that we believe the denoiser prior is expressive enough to capture.
@@
In addition to the average performance, we also compare the performance of these methods on individual images in the right panel of Figure 3A, in a scatter plot of the reconstruction errors from measurements using PCs versus those obtained using OLMs. In both cases, nonlinear reconstruction was performed using the denoiser prior; thus the plot highlights the differences due to the linear measurements. Importantly, we observe improvement in performance for almost all images in the test set for the OLM.
@@ 3.3 Optimized measurements for CelebA
To test the generality of our method for different image classes, we repeated our experiments on the CelebA dataset (Liu et al., 2015), which consists of about $200,000$ centered face images. We resized all images to $80\times 80$ and converted them to grayscale. The rest of the procedure is the same as described above. The bottom half of Figure 2 shows an example face image from the test set. Similar to the MNIST case, we observe that using the denoiser prior to reconstruct from PCs significantly improves the linear inverse estimates. Notably, the conditional samples all appear to be realistic face images even for low $k$; this indicates that the denoiser prior adequately captured the statistical structure of this image dataset. As $k$ increases, the reconstructed face images increasingly resembles the original. Optimizing the measurement matrix offers further improvements in the results, in this case most visible for $k=16$. See Supplemental Figure 4 for two more examples, including reconstructions obtained with random measurements.
The effect of the optimized measurements is quantified in Figure 3B. As with the MNIST dataset, we observe that OLM leads to superior reconstruction using the denoiser prior, illustrated by the red curve in the left panel. The improvements arising from OLM measurements are consistent across all values of $k$, although they are smaller than for the MNIST case. The peak difference in this case is 0.44 dB in PSNR at $k=2048$. Thus, the performance gain between PCA and OLM also depends on the specific image dataset.
Consistent with results on MNIST, reconstruction from PCs using the denoiser prior results in higher performance than linear reconstruction, and all of these methods outperform reconstruction from random measurements. We confirmed that the improvements from the OLM are nearly universal for all images in our test set (Figure 3B, right panel). For completeness, we also used independent component analysis (ICA, Hyvärinen & Oja 2000) to generate linear measurement vectors. On both datasets, we found that reconstructions using the denoiser prior based on ICA measurements were marginally worse than those using the PCs (not shown).

### v4

@@ 2.1 Linear inverse problem
We take a Bayesian statistical approach to solving the inverse problem, in which a prior distribution of the signal, $p(x)$, characterizes the statistical regularities of $x$. Given a partial observation $m$, one can obtain a posterior distribution, $p(x|m)$, and the inverse problem is formulated to minimize an expected loss over this posterior. For squared error loss, the solution is the conditional mean of the posterior, $\hat{x}(m)=\int xp(x|m)\ dx$, and for a “0-1” loss, it is the mode, $\hat{x}(m)=\arg\max_{x}\ p(x|m)$. These solutions are known as minimum mean squared error (MMSE) and maximum a posteriori (MAP) estimates, respectively. More recently, stochastic sampling approaches for solving inverse problems have emerged, where the reconstruction is not the mean or maximum of the posterior, but a high-probability sample (Kadkhodaie & Simoncelli, 2020; Kawar et al., 2022; Chung et al., 2022). We describe the stochastic solution in more detail next.
@@ 2.2 Image prior embedded in a denoiser
where $y$ is the noise-corrupted signal: $y=x+z$, $z\sim\mathcal{N}(0,\sigma^{2}\text{I})$, and $\hat{x}(y)$ is the MMSE denoising solution. This remarkable equation provides an explicit connection between denoising and the density of the noisy image $p_{\sigma}(y)$. See Appendix˜A for the proof of Equation˜1. The distribution of noisy images, $p_{\sigma}(y)$ is related to the image prior $p(x)$ through marginalization:
$p_{\sigma}(y)=\int p(y|x)\,p(x)\,dx=\int g_{\sigma}(y-x)\,p(x)\,dx,$ (2)
where $g_{\sigma}(z)$ is the distribution of Gaussian noise with variance $\sigma^{2}$. This equation expresses the convolution of $p(x)$ with the Gaussian probability density function. That is, $p_{\sigma}(y)$ is a blurred version of $p(x)$ where the extent of blur depends on $\sigma$. The family of $p_{\sigma}(y)$ over all $\sigma>0$ forms a scale-space representation of $p(x)$ and is akin to the temporal evolution of a diffusion process of $p(x)$. A learned denoiser trained over a wide range of $\sigma$ approximates this family of gradients of log densities and can then be used in a coarse-to-fine gradient ascent algorithm to sample from $p(x)$ (see Algorithm˜1).
@@ 2.3 Inverse problem as constrained sampling
See Algorithm˜2 for a detailed description of the constrained sampling algorithm.
@@
which can be approximated by averaging over multiple conditional samples for a given measurement model $M_{k}$, see Section˜E.3 for more details. Visually, however, individual samples will look sharper and of higher visual quality compared to the MMSE estimate (Kadkhodaie & Simoncelli, 2021; Kawar et al., 2022).
@@ 2.4 Optimized linear measurement
In order to solve the optimization problem of Equation˜6, we use stochastic gradient descent in the space of orthonormal matrices. Specifically, we parameterize the set of matrices $Q\in\mathcal{R}^{d\times k}$ with orthonormal columns using the Householder product, which represents matrices as a sequence of elementary reflections as the following (Trefethen & Bau, 2022; Shepard et al., 2015):
@@
[…]
We re-write our empirical objective function from Equation˜5 using the parameterization $Q(\phi)$:
[…]
We search for the optimal measurement matrix within $Q(\phi)$ through variants of gradient descent (Adam optimizer (Kingma & Ba, 2014), see Supplementary Appendix˜E):
@@
In practice, the gradient $\nabla_{\phi_{t}}\mathcal{L}(\phi_{t})$ is approximated by computing the MSE (Equation˜8) on a subset of images sampled from the training set on each iteration. The gradient descent formulation is general, and applicable to any differentiable objective. As a demonstration of this, we also explore a perceptual loss function, the structural similarity index measure (SSIM) (Wang et al., 2004), using the implementation in Detlefsen et al. (2022). See Appendix˜E for details of the datasets, network training, and linear measurement optimization.
@@ 3.1 Two-dimensional example
Figure 1: Reconstruction of 2-D signals from a 1-D linear measurement. Top row: Gaussian distribution. Middle row: K-Sparse distribution. Bottom row: A data distribution confined to a curved one-dimensional manifold. A) Samples from the target distribution (blue points). Measurements are obtained via dot product between samples and a unit vector (black vector) parameterized by angle $\theta$. B) Mean squared error (MSE) of sample reconstruction as a function of measurement angle $\theta$, for two estimators: linear projection onto the measurement vector (blue), and reconstruction using the denoiser prior (orange). The dashed vertical line indicates the optimal solution, $\theta^{*}$, obtained with our optimization procedure (based on Equation˜8). C) Test data drawn from the prior (blue), plotted in a new coordinate system $(x_{1}^{*},x_{2}^{*})$ rotated by $\theta^{*}$ (i.e., such that the optimal measurement $m$ corresponds to the horizontal coordinate $x_{1}^{*}$). Example bivariate estimates sampled from the denoiser prior conditioned on the optimal linear measurement are plotted in red. The MMSE solution (orange) is obtained by averaging over multiple conditional samples.
@@
As an illustration of our method, consider a 1-D measurement problem in a 2-D signal space, for three different prior probabilities (Figure˜1A): a correlated Gaussian distribution, a K-sparse (union of subspace) model, and a tight distribution along a closed 1D manifold. In each case, we train a small, two-layer fully connected denoising network on the data distribution, which is then used to solve the inverse problem via Algorithm˜2. The mean squared reconstruction error obtained by this method as a function of the measurement vector angle is shown in Figure˜1B (red). As a baseline, we compare this to the reconstruction obtained by linear projection onto the measurement axis (Figure˜1B, blue). The optimal measurement vector is obtained by evaluating the reconstruction error for a set of densely sampled 2-D unit measurement vectors spanning orientations $\theta\in[-\pi/2,\pi/2]$ (Figure˜1B, orange). For the optimal measurement vector, Figure˜1C shows samples conditioned on each measurement value (red points) and the MMSE solution as a function of the measurement value (orange line).
@@
We first consider a bivariate Gaussian distribution (Figure˜1, top row). In this case, the first PC is the optimal measurement for both reconstruction methods, as expected for a Gaussian prior. The second row of Figure˜1 depicts the result for a union-of-subspace sparse distribution. Here, the first PC is aligned with the horizontal axis. However, the optimal measurement vector for the denoiser prior is dramatically different. The reconstruction error at the optimal $\theta^{*}$ is near zero. This recapitulates the classical compressed sensing result (Donoho, 2006; Tropp, 2006) that for sparse signals, improved estimates can be achieved by making off-axis measurements.
@@
The last row illustrates the scenario of primary interest in this paper: The data distribution lies in a low-dimensional but curved manifold. This type of higher-order structure cannot be adequately captured by either the Gaussian or sparse prior, but can be effectively described using the more powerful diffusion models, such as our denoiser prior (Supplemental Figure˜2). Notably, in this case, we also identified an optimal measurement angle that outperforms the principal axis. Importantly, the optimal angle $\theta^{*}$ is correctly identified by our optimization method across all three cases, illustrating the generality of our approach.
@@
Finally, we highlight another aspect of the reconstruction demonstrated by the 2D examples (Figure˜1C). Our methods generate estimates by sampling from the denoiser prior, conditioned on the linear measurements. While individual samples are consistent with the conditional prior, the MMSE solutions are obtained by averaging multiple conditional samples. This averaging, however, can cause deviations of the MMSE solution from the prior. In the context of image reconstruction, the MMSE solution often exhibits lower visual quality (e.g., blurrier images) compared to individual conditional samples, as it does not lie close to the image prior manifold (Kadkhodaie & Simoncelli, 2021; Kawar et al., 2022). This observation is particularly relevant in the results presented below and further motivates our approach of optimizing measurements with respect to a perceptual loss function later.
@@ 3.2 Optimized measurements for MNIST
We used our method to obtain optimized measurements for the MNIST dataset (Deng, 2012). We first trained a DNN denoiser on these digit images – for details of the architecture and training, see Appendix˜E. We apply our method as described in Section˜2.4 to obtain the optimized measurement matrix (OLM) for this dataset, for a range of $k$ values.
@@
For comparison, we apply our reconstruction method to two other types of linear measurements: the top $k$ PCs, and $k$ random vectors, which are optimal measurements under Gaussian and sparse prior assumptions, respectively. We also compare to the optimal reconstruction for a Gaussian prior, which is simply linear projection onto the space spanned by the top $k$ PCs. The top half of Figure˜2 illustrates our results for a test digit image. First, we observe that combining the PCs with the denoiser prior significantly improves the linear inverse estimates, which reflects the power of the denoiser prior. All conditional samples from the denoiser prior appear to be real MNIST images, which then gradually converge to the original image as $k$ increases. Use of the optimal measurement matrix (OLM) offers additional improvements, reconstructing the correct digit with a smaller $k$. These improvements are evident in both the identity of the digits and their more detailed appearance. Using random measurements (not shown) does not perform as well as either the optimized linear measurements or PCA. See Supplemental Figure˜3 for two more examples, including random measurement vectors.
@@
Figure 3: Performance comparisons for A) MNIST and B) CelebA datasets. Left graph in each panel shows the peak signal-to-noise ratio (PSNR, in dB) as a function of the number of measurements $k$, for the denoiser prior reconstruction from OLMs (red), denoiser prior reconstruction from PCs (orange), linear (projection) reconstruction from PCs (blue), and denoiser prior reconstruction from random measurements (green). Inline numbers indicate the increase in PSNR from PC + denoiser prior to OLM + prior. The right graph in each panel plots the PSNR of the PC reconstruction against that of the OLM reconstruction over all images in the test set, for a value of $k$ approximately $10\%$ of the total number of pixels. Red points indicate the examples shown in Figure˜2 and Supplemental Figure˜3 and Figure˜4. Also see Supplemental Table˜1 for comparison of our results to other previous methods.
@@
We quantify our results in terms of peak signal-to-noise ratio (PSNR), as a function of $k$ in the left panel of Figure˜3A. The OLMs result in superior reconstruction (red curve), particularly for the values of $k$ ranging from $1\%$ to $30\%$ of the total number of pixels. The peak performance gain of OLM over PC (both utilizing denoiser prior) reached 4.72 dB in PSNR at $k=256$. Between the two PC reconstructions, the nonlinear prior-based method (orange) shows a considerable improvement over linear reconstruction (blue), demonstrating the substantial advantage of using the more complex denoiser prior. See Supplemental Table˜1 for a comparison of our results to other methods in the literature. Lastly, consistent with previous reports using simpler image priors (Weiss et al., 2007), we observe that PC measurements outperform random measurements when used with the denoiser prior (green). This raises the possibility that the union of subspace prior assumed in the compressed sensing literature (Donoho, 2006) does not fully capture the structure of these images, since random measurements are expected to perform well if the images indeed lie on a union of subspaces – a structure that we believe the denoiser prior is expressive enough to capture.
@@
In addition to the average performance, we also compare the performance of these methods on individual images in the right panel of Figure˜3A, in a scatter plot of the reconstruction errors from measurements using PCs versus those obtained using OLMs. In both cases, nonlinear reconstruction was performed using the denoiser prior; thus the plot highlights the differences due to the linear measurements. Importantly, we observe improvement in performance for almost all images in the test set for the OLM.
@@ 3.3 Optimized measurements for CelebA
To test the generality of our method for different image classes, we repeated our experiments on the CelebA dataset (Liu et al., 2015), which consists of about $200,000$ centered face images. We resized all images to $80\times 80$ and converted them to grayscale. The rest of the procedure is the same as described above. The bottom half of Figure˜2 shows an example face image from the test set. Similar to the MNIST case, we observe that using the denoiser prior to reconstruct from PCs significantly improves the linear inverse estimates. Notably, the conditional samples all appear to be realistic face images even for low $k$; this indicates that the denoiser prior adequately captured the statistical structure of this image dataset. As $k$ increases, the reconstructed face images increasingly resembles the original. Optimizing the measurement matrix offers further improvements in the results, in this case most visible for $k=16$. See Supplemental Figure˜4 for two more examples, including reconstructions obtained with random measurements.
The effect of the optimized measurements is quantified in Figure˜3B. As with the MNIST dataset, we observe that OLM leads to superior reconstruction using the denoiser prior, illustrated by the red curve in the left panel. The improvements arising from OLM measurements are consistent across all values of $k$, although they are smaller than for the MNIST case. The peak difference in this case is 0.44 dB in PSNR at $k=2048$. Thus, the performance gain between PCA and OLM also depends on the specific image dataset.
Consistent with results on MNIST, reconstruction from PCs using the denoiser prior results in higher performance than linear reconstruction, and all of these methods outperform reconstruction from random measurements. We confirmed that the improvements from the OLM are nearly universal for all images in our test set (Figure˜3B, right panel). For completeness, we also used independent component analysis (ICA, Hyvärinen & Oja 2000) to generate linear measurement vectors. On both datasets, we found that reconstructions using the denoiser prior based on ICA measurements were marginally worse than those using the PCs (not shown).